OpenStudy (frankypoo):

Find the derivative of the function: √-9-5x. dy/dx=? help much appreciated :)

OpenStudy (calculusfunctions):

Is that\[y =\sqrt{-9-5x}\]

Parth (parthkohli):

You can use the Chain Rule. I'm sure Sir @calculusfunctions will tell you more.

OpenStudy (frankypoo):

@calculusfunctions, yeah

Parth (parthkohli):

\[\rm {d \over dx}f(g(x)) = g'(x)f'(g(x))\]

OpenStudy (calculusfunctions):

OK then as @ParthKohli said very accurately, do you know the chain rule or the power of a function rule?

OpenStudy (frankypoo):

^ lol

OpenStudy (frankypoo):

@ParthKohli got it

Parth (parthkohli):

Yes, that's another one you can use! Though you have to use Chain Rule, and the power rule comes in that.

OpenStudy (calculusfunctions):

Yes @ParthKohli that is excellent by there is an easier rule which is also a form of the chain rule.

OpenStudy (calculusfunctions):

@frankypoo Give me a few minutes to type up a explanation for you.

OpenStudy (frankypoo):

Alright (:

Parth (parthkohli):

He's a great teacher, isn't he?

OpenStudy (frankypoo):

He knows what he's doing aha, I'm doing calc atm, It's all in his name aha

OpenStudy (frankypoo):

doing Derivatives and with all of the rules and types of problems I get mixed up :c

OpenStudy (calculusfunctions):

Chain Rule (Power of a Function Rule):\[y =C[(f(x)]^{n}\]where C and n are any real number. \[\frac{ dy }{ dx }=Cn[f(x)]^{n -1}f \prime(x)\] Example:\[y =5\sqrt[5]{x ^{2}+2}\]First rewrite\[y =5(x ^{2}+2)^{\frac{ 1 }{ 5 }}\]Now\[\frac{ dy }{ dx }=(5)(\frac{ 1 }{ 5 })(x ^{2}+2)^{\frac{ 1 }{ 5 }-1}(\frac{ d(x ^{2}+2) }{ dx })\] \[\frac{ dy }{ dx }=(x ^{2}+2)^{\frac{ -1 }{ 4 }}(2x)\] \[\frac{ dy }{ dx }=\frac{ 2x }{ \sqrt[4]{x ^{2}+2} }\]

OpenStudy (calculusfunctions):

@frankypoo if you understand then try your question now. @ParthKohli thank you and have you seen this form of the chain rule before? It is still the Chain rule but it's called the Power of a function rule.

Parth (parthkohli):

No, I've not... I'm currently learning about it. Thank you sir!

OpenStudy (calculusfunctions):

Of course @ParthKohli Nice profile picture.

Parth (parthkohli):

Thanks sir, and yours too! (I love Sheldon Cooper :) )

OpenStudy (calculusfunctions):

@frankypoo where are you?

OpenStudy (frankypoo):

here ahaha.

OpenStudy (frankypoo):

"scooby doo, where are you?!" just reminded me of that

OpenStudy (frankypoo):

I'm about to try it, give me a sec.

OpenStudy (calculusfunctions):

Yes lol

OpenStudy (calculusfunctions):

Take your time. I'm just humming the theme song, accept I replace Scooby Doo with @frankypoo in the lyrics. LOL

OpenStudy (frankypoo):

ahaha, alright

Parth (parthkohli):

lol

OpenStudy (frankypoo):

I'm still confused :s

OpenStudy (calculusfunctions):

With which part?

OpenStudy (frankypoo):

EVERYTHING

OpenStudy (frankypoo):

LOL

OpenStudy (frankypoo):

I think I need another example.

OpenStudy (calculusfunctions):

OK we'll take it one step at a time. Babypoo steps! LOL but first I need to know if you at least understand the the rule? If not that's OK too because you will once I'm through.

OpenStudy (frankypoo):

unless the answer is... \[\frac{ dy }{ dx }=\frac{ x }{ \sqrt{-5x-9}}\]

OpenStudy (calculusfunctions):

NO!

OpenStudy (frankypoo):

which I doubt

OpenStudy (frankypoo):

LOL

OpenStudy (calculusfunctions):

OK step 1: Can you first rewrite the function without radicals exactly like I did in the example?

OpenStudy (frankypoo):

y=(-9-5x)

OpenStudy (calculusfunctions):

No...What are you missing?

OpenStudy (frankypoo):

.-.

OpenStudy (calculusfunctions):

\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]using this fact, try again!

OpenStudy (frankypoo):

dafuq

OpenStudy (calculusfunctions):

Do you know the rational exponent law?

OpenStudy (frankypoo):

noo,

OpenStudy (frankypoo):

ill search it real quick.

OpenStudy (calculusfunctions):

No, I'll teach it to you @frankypoo

OpenStudy (frankypoo):

Teach it quick-like

OpenStudy (calculusfunctions):

A rational power, m/n, raised to the base x is equal to the nth root of x raised to the mth power or the mth power of the nth root of x. In other words \[x ^{\frac{ m }{ n }}=(\sqrt[n]{x})^{m}=\sqrt[n]{x ^{m}}\] Example:\[8^{\frac{ 2 }{ 3 }}=(\sqrt[3]{8})^{2}=2^{2}=4\]

OpenStudy (frankypoo):

makes sense

OpenStudy (calculusfunctions):

Please do not rush me when I am teaching in this format. I type slowly with care to make sure that no errors are made in the lessons I give you. Alright?

OpenStudy (calculusfunctions):

OK so then if it makes sense then can you rewrite your function without radicals again?

OpenStudy (calculusfunctions):

@frankypoo I had to go half an hour earlier but I want to finish teaching you first. My point is that I didn't rush you to learn faster so then perhaps you should extend the same courtesy.

OpenStudy (calculusfunctions):

Do you need an example of how to use the rational exponent law to rewrite the function without radicals?

OpenStudy (frankypoo):

would be appreciated

OpenStudy (calculusfunctions):

Suppose \[y =\sqrt{3x ^{5}-x +4}\] Then by the rational exponent law \[y =(3x ^{5}-x +4)^{\frac{ 1 }{ 2 }}\] Understand? Now can you rewrite yours?

OpenStudy (frankypoo):

where did the 1/2 come from?

OpenStudy (calculusfunctions):

I thought you said you understood the rational power law. Please take another look while I explain again.

OpenStudy (frankypoo):

y=(-5x-9)^1/2

OpenStudy (calculusfunctions):

Excellent!!! Understand now??

OpenStudy (calculusfunctions):

If you have a nth root of a function f(x) raised to the exponent m, then the function f(x) is raised to the exponent m/n.

OpenStudy (frankypoo):

yea, bit more sense

OpenStudy (calculusfunctions):

\[y =(-9-5x)^{\frac{ 1 }{ 2 }}\]According to what I demonstrated, what is f(x)?

OpenStudy (calculusfunctions):

What is the exponent and what is the number in front of the bracket?

OpenStudy (calculusfunctions):

Can you please tell me at least that much?

OpenStudy (frankypoo):

I'm trying

OpenStudy (calculusfunctions):

Baby steps remember? All I asked is if\[y =(-9-5x)^{\frac{ 1 }{ 2 }}\]then what is the number in front of the brackets and what is the exponent?

OpenStudy (calculusfunctions):

There is no number visible in front of the bracket so it's a 1. What is the exponent?

OpenStudy (frankypoo):

in front of the brackets there's an invisible 1, exponent 1/2

OpenStudy (calculusfunctions):

Very good! Now you multiply those numbers so what do you get?

OpenStudy (calculusfunctions):

What is 1 multiplied by 1/2?

OpenStudy (frankypoo):

1/2.

OpenStudy (calculusfunctions):

Excellent! Now if you raise the function to an exponent one less than what it is now, what do you get?

OpenStudy (calculusfunctions):

(1/2) - 1 = ?

OpenStudy (frankypoo):

-1/2

OpenStudy (calculusfunctions):

Correct! Now finally if you take the derivative of the function inside the bracket, what do you get?

OpenStudy (frankypoo):

-5

OpenStudy (calculusfunctions):

What is the derivative, dy/dx, if y = -9 - 5x ?

OpenStudy (calculusfunctions):

Right!

OpenStudy (frankypoo):

\[\frac{ -5 }{ \sqrt{-5x-9} }\]

OpenStudy (frankypoo):

what am I missing?

OpenStudy (calculusfunctions):

Almost! What's missing? Where is the 1/2?

OpenStudy (frankypoo):

isnt it 2 by the radical?

OpenStudy (calculusfunctions):

And please write it properly as dy/dx = ?

OpenStudy (calculusfunctions):

That's what I meant by where is the 1/2 ?

OpenStudy (calculusfunctions):

So can you please write your final answer again with dy/dx?

OpenStudy (frankypoo):

\[\frac{ dy }{ dx }=\frac{ -5 }{ \sqrt[2]{-5x-9} }\]

OpenStudy (calculusfunctions):

NO! It should look like this\[\frac{ dy }{ dx }=\frac{ -5 }{ 2\sqrt{-5x -9} }\]You have your 2 in the wrong spot!!

OpenStudy (calculusfunctions):

Understand now?

OpenStudy (frankypoo):

I meant to do that :D

OpenStudy (frankypoo):

yessir, thanks for your help and your patience.

OpenStudy (calculusfunctions):

Very Good! @frankypoo congratulations on finally solving another mystery just like Scooby Doo.

OpenStudy (frankypoo):

hah, alright!:)

OpenStudy (calculusfunctions):

You're welcome! Anytime.

OpenStudy (calculusfunctions):

I don't even get a medal? LOL

OpenStudy (frankypoo):

there ya go :)